OpenStudy (anonymous):

Help how do I do this?? find domain for f+g(x) f(x)=1/x-3 g(x)=2x+4

OpenStudy (australopithecus):

is that f(x) + g(x)? Can you show me what f(x) + g(x) equals

OpenStudy (australopithecus):

or rather express it in a way in which we can find its domain

OpenStudy (anonymous):

@Australopithecus it just asked for the domain of f(x) + g(x) and all that is given is f(x)=1/x-3 g(x)=2x+4

OpenStudy (anonymous):

you solve f and then you solve g and then you get the solutions and then you get ur answer

OpenStudy (anonymous):

so the solution is the domain? @peanutbutter817

OpenStudy (australopithecus):

first off you need to understand function notation. f(x) is just another way of writing y, but it provides more information because you can use it to designate a point in a function readily by simply replacing x with a number for example: y = x + 3 can be represented as, f(x) = x + 3 when we say f(3) we simply mean f(3) = 3 + 3 just realize that, f(x) is the same as writing x + 3 so if you have, f(x) + g(x) where, f(x) = x + 3 and g(x) = 2 + x f(x) + g(x) = (X + 3) + (2+x)

OpenStudy (anonymous):

once you solve each g and f plug them in f(x)+g(x)and then you will get your answer @eragon4

OpenStudy (australopithecus):

once you write out what f(x) + g(x) is in the case of your problem you can look for the domain, the domain is simply the numbers that can be inputed as x in your equation. You have to remember simple rules to recognize what numbers cannot be inputed into your function (and thus figure out what your domain is). Rules: 1. You cannot divide by 0, meaning zero cannot be in the denominator, it can be in the numerator though, in which case the fraction always equals 0) To clarify: 3/0 = undefined 3/(3-3) = undefined 0/0 = undefined 0/3 = 0 2. In the case of radicas (numbers raised to the power of a fraction) you cannot have a negative number under them you can have 0 though, as this will give you imaginary numbers (The domain is simply concerned with what real numbers are allowed). To clarify: \[\sqrt{-1} = Imaginary Number\] \[-3^\frac{1}{3} = Imaginary Number\] etc. 3. When dealing with logrithems such as log(x) or ln(x), x cannot be less than or equal to 0 To clarify: Recognize that log(x) = y is the same as 10^y = x in this case you should be able to recognize that in 10^y = x, there is no number you can input into y that will make x equal to or less than zero

OpenStudy (australopithecus):

I recommend studying what I wrote above then we can go over the problem together, all I ask is you express, f(x) + g(x) in the way I specified above